Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

PS - Probability - 2 - exactly one

Expert replies
by karthikpandian19 » Tue Jul 24, 2012 5:10 am
Two of six students in a seminar are to be randomly assigned to read On Liberty. If Deringer and Hay are both in the section, what is the probability that exactly one of them is assigned to read On Liberty?


(A) 4/15

(B) 2/5

(C) 8/15

(D) 3/5

(E) 2/3
Regards,
Karthik
The source of the questions that i post from JUNE 2013 is from KNEWTON

---If you find my post useful, click "Thank" :) :)---
---Never stop until cracking GMAT---
Join the discussion
Source: — Problem Solving |

by niketdoshi123 » Tue Jul 24, 2012 5:49 am
karthikpandian19 wrote:Two of six students in a seminar are to be randomly assigned to read On Liberty. If Deringer and Hay are both in the section, what is the probability that exactly one of them is assigned to read On Liberty?


(A) 4/15

(B) 2/5

(C) 8/15

(D) 3/5

(E) 2/3
Total possible selections = 6C2 = 15
Selecting exactly one out of Deringer and Hay = 2 ways
Selecting others = 4 ways
Total ways = 2*4 = 8
Probability = [spoiler]8/15[/spoiler]
Ans is c
Join the discussion

by CSASHISHPANDAY » Wed Jul 25, 2012 5:38 am
Let pick D AND not H:

(2/6)*(4/5)= 8/30



same way pick H AND not picking D:

(2/6)*(4/5)= 8/30



Probability of picking only one of them: sum both probabilities:

(8/30)+(8/30)=8/15


Let's say you're interested in picking D then picking Not H. The probability of picking D on the first draw is 1/6. Then there are 5 people left out of the original 6. Of those 5 people, you are interested in 4 of them, because the remaining person is H, and you don't want to pick him. Thus, the probability of picking Not H on the second draw is 4/5.
Join the discussion

by NicoleWhite » Wed Jul 25, 2012 1:53 pm
I am confused as to why the probability of picking D is 2/6.
Let pick D AND not H:

(2/6)*(4/5)= 8/30
D is 1 person among 6 people, so why are you using 2/6?

When I solve this, I do P(D)*P(not H) + P(H)*P(not D), which is (1/6)*(4/5) + (1/6)*(4/5) = (4/30) + (4/30) = (8/30) = (4/15)

Can you explain where I am going wrong?
Join the discussion

by eagleeye » Wed Jul 25, 2012 2:11 pm
NicoleWhite wrote:I am confused as to why the probability of picking D is 2/6.
Let pick D AND not H:

(2/6)*(4/5)= 8/30
D is 1 person among 6 people, so why are you using 2/6?

When I solve this, I do P(D)*P(not H) + P(H)*P(not D), which is (1/6)*(4/5) + (1/6)*(4/5) = (4/30) + (4/30) = (8/30) = (4/15)

Can you explain where I am going wrong?
I can't comment on the thought process CSASHISHPANDAY used to get 2/6, but I can tell you where you are going wrong.

The way you are calculating it is equivalent to doing this:

1. Pick D first. Pick one of the people other than H second.
2. Pick H first. Pick one of the people other than D second.

This is a permutational way of calculating probability. You are missing two cases. Those are:
1. Pick one of people other than D and H first. Pick D second.
2. Pick one of people other than D and H first. Pick H second.

This would give you the correct probability of 8/15.

Another way you could do it using 2/6 is:

1. (Pick either one of D or H first) * (Pick one of the other two people second)*2!
= 2/6*4/5*2!.
The 2! above is for arranging/considering the case where one of D or H is picked second.


The fastest way using selections of doing this question would be:

Required probability = (Choose one of D or H)*(Choose one of the other 4)/(Select 2 out of 6)
= 2C1*4C1/6C2 = 2*4*2/(6*5) = 8/15.

Let me know if this helps :)
Join the discussion

by NicoleWhite » Wed Jul 25, 2012 3:40 pm
Ahh, I did not account for the cases you described:
1. Pick one of people other than D and H first. Pick D second.
2. Pick one of people other than D and H first. Pick H second.
If I continued with my method, I would get P(not D or H)*P(D) = (4/6)*(1/5) = (4/30) and P(not D or H)*P(H) = (4/6)*(1/5) = (4/30)

Adding these to the other (4/30)'s would get me to the correct answer. Thank you!
Join the discussion

by GMATGuruNY » Wed Jul 25, 2012 4:52 pm
karthikpandian19 wrote:Two of six students in a seminar are to be randomly assigned to read On Liberty. If Deringer and Hay are both in the section, what is the probability that exactly one of them is assigned to read On Liberty?


(A) 4/15

(B) 2/5

(C) 8/15

(D) 3/5

(E) 2/3
P(D or H on the first pick) = 2/6. (Of the 6 students, 2 are D or H).
P(not D or H on the second pick) = 4/5. (Of the 5 remaining students, 1 is D or H, so the remaining 4 are not D or H).
Since we want both events to happen, we multiply the fractions:
2/6 * 4/5.
Since the order the picks could be reversed to not D or H on the first pick followed by D or H on the second pick, we multiply by 2:
2/6 * 4/5 * 2 = 8/15.

The correct answer is C.

Alternate approach:
P(D and H) = 2/6 * 1/5 = 1/15.
P(neither D nor H) = 4/6 * 3/5 = 6/15.
P(D or H but not both) = 1 - 1/15 - 6/15 = 8/15.
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion

by CSASHISHPANDAY » Wed Jul 25, 2012 8:01 pm
Dear nicokle, read once again

Two of six students in a seminar are to be randomly assigned to read On Liberty. If Deringer and Hay are both in the section, what is the probability that exactly one of them is assigned to read On Liberty?

Here you have to pick two students which may be either D or H (two person) either way like D and not H/H but not D so pick should be 2/6
let it be simpler P(D or H be the 1st pick i.e. out of six 2 D or H)=2/6
Join the discussion

by karthikpandian19 » Wed Aug 01, 2012 7:14 pm
OA is C.....thanks for all of your explanations guys...
Regards,
Karthik
The source of the questions that i post from JUNE 2013 is from KNEWTON

---If you find my post useful, click "Thank" :) :)---
---Never stop until cracking GMAT---
Join the discussion